OpenStudy (anonymous):

Evaluate the triple integral ∫∫∫ xy dV where E is bounded by the parabolic cylinders y =x^2 and x = y^2 and the planes z = 0 and z = x + y.

OpenStudy (anonymous):

I can't draw out the bounded region for parabolic cylinders y =x^2 and x=y^2 can someone help me find the limits

OpenStudy (anonymous):